### v1

@@ F.1 The Details of Cloud Guessing Game (CGG)
Specifically, the data construction process of CGG is as follows: First, we consider four categories—cat, human, and giraffe. We selected unambiguous and categorically distinct images for each category from the Internet. Taking the example of a cat, we chose the image shown in Fig.9 (a) in main text, then use Photoshop carefully to create a mask of the cat as shown in Fig.9 (b) in main text. Subsequently, employing the control diffusion model [31, 32, 84, 85] with the following prompt, we generated a cat-like white cloud as depicted in Fig.9 (c) in main text. (Prompt for white cloud generation) masterpiece, best quality, white cloud++, Stratus cloud,Altostratus cloud,Cirrus cloud, blue sky,light rays, sharp focus, HDR, UHD, 8K, masterpiece, Highly detailed, extreme detail detail, reality,realistic light, real, physics, reality, photo reality, Deconstruction Continuously generated through the control diffusion model, we manually screen until we identify 30 unambiguous and difficulty challenging white cloud images for each category. The difficulty is adjusted by the “controlnet_scale,” a coefficient used to control the intensity of mask control. A higher value implies a stronger correlation between the generated images and masks, resulting in lower difficulty. Finally, employing the instruction template from Fig.6 (a) in the main text, we construct choice questions for 4T1, with options randomly arranged from both the ground truth category and three words sampled randomly from the unrelated word set [‘chair’, ‘cup’, ‘sing’, ‘jump’, ‘rap’, ‘basketball’, ‘computer’, ‘egg’, ‘phone’, ‘house’, ‘lamp’, ‘shoes’], ensuring each question’s validity and clarity through manual verification.
@@ F.2 The Details of Divergent Association Task (DAT)
Please carefully understand the provided question and select the option that satisfies the problem. Only one option meets the requirements. Question: Please select the option least relevant to the current set of words. Words: <Words> Options: <Options> Answer Format: Please respond in the format of ’Option id. Option content,’ for example, ’A. xxx.’ Response: Satisfactory option is <Response>
@@
Example 1: <Words>: Guitar Amplifier Strings Pick Melody Chord Song Musician Concert <Options>: A.studio B.hat C.piano D.umbrella
Example 2: <Words>: Guitar Amplifier Strings Pick Melody Chord Song Musician Concert <Options>: A.flame B.orange C.diamond D.earth
Example 3: <Words>: Soccer Amplifier Marathon Surfing Volleyball Basketball Carrot Running Yoga <Options>: A.canvas B.wire C.volcano D.bracelet
Example 4: <Words>: Pepper Zucchini Eggplant Surfing Garlic Potato Carrot Bean Gymnastics <Options>: A.drill B.bee C.hourglass D.brick
Example 5: <Words>: Decaf Pastry Brew Roast Forest Outdoors Compass Bean Backpack <Options>: A.cake B.whip C.space D.river

### v3

@@ F.1 The Details of Cloud Guessing Game (CGG)
Specifically, the data construction process of CGG is as follows: First, we consider four categories—cat, human, and giraffe. We selected unambiguous and categorically distinct images for each category from the Internet. Taking the example of a cat, we chose the image shown in Fig.9 (a) in main text, then use Photoshop carefully to create a mask of the cat as shown in Fig.9 (b) in main text. Subsequently, employing the control diffusion model [31, 32, 85, 86] with the following prompt, we generated a cat-like white cloud as depicted in Fig.9 (c) in main text. {mdframed}[backgroundcolor=gray!8] (Prompt for white cloud generation) masterpiece, best quality, white cloud++, Stratus cloud,Altostratus cloud,Cirrus cloud, blue sky,light rays, sharp focus, HDR, UHD, 8K, masterpiece, Highly detailed, extreme detail detail, reality,realistic light, real, physics, reality, photo reality, Deconstruction Continuously generated through the control diffusion model, we manually screen until we identify 30 unambiguous and difficulty challenging white cloud images for each category. The difficulty is adjusted by the “controlnet_scale,” a coefficient used to control the intensity of mask control. A higher value implies a stronger correlation between the generated images and masks, resulting in lower difficulty. Finally, employing the instruction template from Fig.6 (a) in the main text, we construct choice questions for 4T1, with options randomly arranged from both the ground truth category and three words sampled randomly from the unrelated word set [‘chair’, ‘cup’, ‘sing’, ‘jump’, ‘rap’, ‘basketball’, ‘computer’, ‘egg’, ‘phone’, ‘house’, ‘lamp’, ‘shoes’], ensuring each question’s validity and clarity through manual verification.
@@ F.2 The Details of Divergent Association Task (DAT)
{mdframed}
[backgroundcolor=gray!8] Please carefully understand the provided question and select the option that satisfies the problem. Only one option meets the requirements. Question: Please select the option least relevant to the current set of words. Words: <Words> Options: <Options> Answer Format: Please respond in the format of ’Option id. Option content,’ for example, ’A. xxx.’ Response: Satisfactory option is <Response>
@@
{mdframed}
[backgroundcolor=gray!8] Example 1: <Words>: Guitar Amplifier Strings Pick Melody Chord Song Musician Concert <Options>: A.studio B.hat C.piano D.umbrella
{mdframed}
[backgroundcolor=gray!8] Example 2: <Words>: Guitar Amplifier Strings Pick Melody Chord Song Musician Concert <Options>: A.flame B.orange C.diamond D.earth
{mdframed}
[backgroundcolor=gray!8] Example 3: <Words>: Soccer Amplifier Marathon Surfing Volleyball Basketball Carrot Running Yoga <Options>: A.canvas B.wire C.volcano D.bracelet
{mdframed}
[backgroundcolor=gray!8] Example 4: <Words>: Pepper Zucchini Eggplant Surfing Garlic Potato Carrot Bean Gymnastics <Options>: A.drill B.bee C.hourglass D.brick
{mdframed}
[backgroundcolor=gray!8] Example 5: <Words>: Decaf Pastry Brew Roast Forest Outdoors Compass Bean Backpack <Options>: A.cake B.whip C.space D.river